OpenStudy (anonymous):

Find the exact area of the surface obtained by rotating the curve about the x-axis. y=sqrt(1+4x) 1

OpenStudy (anonymous):

Are you allowed to use Calculus?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

That's what I'm trying to do. I got up to \[2\pi \int\limits_{1}^{5} \sqrt{1+4x} \sqrt{1/\sqrt{1+4x}}\]

OpenStudy (anonymous):

That formula looks a bit confusing to me CatLove9, try to think what the solid looks like, the radius especially.

OpenStudy (dumbcow):

the 2nd radical isn't quite right for surface area integral, multiply by dA \[dA = \sqrt{1+f'(x)^{2}}\] \[f'(x) = \frac{2}{\sqrt{4x+1}}\] \[\rightarrow \sqrt{1+\frac{4}{4x+1}} = \sqrt{\frac{4x+5}{4x+1}}\]

OpenStudy (anonymous):

thought it would be \[1\over2\sqrt{4x+1}\]

OpenStudy (dumbcow):

chain rule...multiply that by derivative of inside

OpenStudy (anonymous):

Oh, yes yes, I forgot about that.

OpenStudy (anonymous):

How did you get the 4x on the top again? I forgot how to do that.

OpenStudy (dumbcow):

combining fractions change the 1 to 4x+1/4x+1

OpenStudy (anonymous):

Okay I got you. so then what happens after that.

OpenStudy (dumbcow):

you have to integrate...it comes out nicely because the sqrt(1+4x) cancels out \[2\pi \int\limits_{1}^{5}\sqrt{1+4x}*\frac{\sqrt{4x+5}}{\sqrt{1+4x}} dx\]

OpenStudy (anonymous):

I see, thank you!

OpenStudy (dumbcow):

your welcome...for an answer i got 98pi/3